### v1

# Advances in Robust Federated Learning: Heterogeneity Considerations

Abstract
In the field of heterogeneous federated learning (FL), the key challenge is to efficiently and collaboratively train models across multiple clients with different data distributions, model structures, task objectives, computational capabilities, and communication resources. This diversity leads to significant heterogeneity, which increases the complexity of model training. In this paper, we first outline the basic concepts of heterogeneous federated learning and summarize the research challenges in federated learning in terms of five aspects: data, model, task, device and communication. In addition, we explore how existing state-of-the-art approaches cope with the heterogeneity of federated learning, and categorize and review these approaches at three different levels: data-level, model-level, and architecture-level. Subsequently, the paper extensively discusses privacy-preserving strategies in heterogeneous federated learning environments. Finally, the paper discusses current open issues and directions for future research, aiming to promote the further development of heterogeneous federated learning.
Index Terms:
Federated Learning, heterogeneity, robust, privacy-preserving
I Introduction
With the popularization of mobile devices and 5G networks, a large amount of data is generated on edge devices every day, greatly promoting the development of artificial intelligence [1]. Traditional machine learning (ML) methods require centralized data storage and processing for model training, which poses significant challenge to data transmission and storage. Moreover, with the frequent occurrence of data leakage, the issue of privacy protection attracts more and more attention. In this context, institutions and enterprises are going to seek a new ML paradigm, which can utilize distributed data resources for training while protecting data privacy and security.
In 2016, Google researchers proposed the concept of federated learning (FL) [2]. The core idea about FL is to train and update models locally on data owners (i.e. clients) and aggregates model updates on a third party (i.e. server), without the need to transfer the raw data to a central server. By this way, FL can significantly reduce data transmission in the network, thus reducing the possibility of privacy leakage and the requirement for centralized computing and storage resources. FL not only presents the respect and protection for privacy, but also demonstrates how to maximize the utility of data while ensuring privacy in the era of big data. At present, FL is becoming an effective way for balancing the utility and privacy of data, bringing new development directions and opportunities to ML. FL has demonstrated its broad application potential in fields such as intelligent healthcare [3] finance [4], the Internet of Things [5], and smart cities [6].
However, in the application scenarios of typical FL, there are multiple heterogeneity issues among clients, including data heterogeneity, model heterogeneity, task heterogeneity, communication heterogeneity, and device heterogeneity, which brings challenges to collaboratively model training. The presence of heterogeneity not only significantly damages the generalization ability and prediction accuracy of trained model, but also to some extent restrict the application of FL. Specifically, FL often counters the following heterogeneity: Data heterogeneity: The clients often far away from each other geographically and lack communications, and each client may collect data from different data source, causing skews in the distributions of data collected by clients [7]. Moreover, the method of sampling and processing data and the amount of collected data may also vary across different clients, resulting in quality skew and quantity skew. Model heterogeneity: Due to device limitations, task types and other reasons, the clients have to customize their own model structures and jointly train models in a federated manner, which is model structure heterogeneity [8]. Moreover, the adopted optimizer of different clients may also be different. Task heterogeneity: Sometimes different clients need to train models for different tasks. For example, some clients may need to train models for image classification, while some other clients may need to train models for semantic segmentation.[9]. Communication heterogeneity: In typical FL, each round of FL requires transmission of model parameters or gradients between the server and the clients. But in reality, the participants of FL are usually under different network environments, and communication bandwidth and budgets of some participants are limited, which is called communication heterogeneity [10]. Device heterogeneity: The typical FL algorithms usually require local training on the clients. But due to limited computing power, storage resources, energy supply, etc., the clients may update local models inconsistently and thus lower the efficiency of the whole system [11].
Although FL is originally proposed for helping protect privacy of the clients, the risk of privacy leakage has not been completely eliminated. Besides heterogeneity, FL is still subject to some privacy attacks. The reason is that attackers can easily access client uploads to analyze privacy information. For example, the malicious attackers may intercept the uploaded parameters or gradients and reconstruct the client’s local data by Gradient Inversion [12] or Model Inversion [13]. Moreover, attackers may also use Membership Inference Attack [14] to determine whether certain data is includes in local training, or use Property Inference Attack to infer private information of some property, like the data sampling environment and label distribution [15]. Thus, it is necessary to employ some privacy-preserving mechanisms in FL to enable clients to communicate with servers and other clients in a more secure way.
Research addressing heterogeneous FL is currently generating a lot of buzz, and we investigate a number of frequently cited or recent surveys related to heterogeneity and privacy issues. Although they have all provided detailed descriptions and abundant related researches for their respective focused topics, they still overlook some less mentioned types of issues and techniques. In this paper, we first categorize several types of heterogeneity in detail, and explore heterogeneity challenges may encountered while applying FL for practical environments, as well as solutions. For each type of heterogeneity, we not only present detailed and formal description, but also conduct extensive survey for comprehensively summarizing the state-of-the-art solutions. Moreover, given that privacy protection is one of the core demands of FL, we also focus on privacy protection problems and solutions while applying FL in scenarios with heterogeneity. We discuss the application of existing privacy protection techniques in FL, delving into the trade-off between privacy and utility, and introduce the advanced researches in relevant fields. We list the differences between our survey and the above other surveys in Table I.
TABLE I: Comparison between other surveys and ours
Survey [16] [17] [18] [19] [11] [20] [21] [22] [23] Ours
Data Heterogeneity Fig. 1 Distribution Skew $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Label Skew $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\checkmark$
Feature Skew $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Quality Skew $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\checkmark$
Quantity Skew $\checkmark$ $\times$ $\times$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ $\checkmark$
Model Heterogeneity Fig. 2 Structure Heterogeneity $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\checkmark$
Optimizer Heterogeneity $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\checkmark$
Task Heterogeneity Fig. 3 $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$
Communication Heterogeneity $\checkmark$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Device Heterogeneity $\checkmark$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Privacy Mechanism Fig. 4 Differential Privacy $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Homomorphic encryption $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Model Watermark $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\checkmark$
Blockchain $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$
Thus, the main contributions of this work are threefold:
  1. 1.
    We systematically categorize and analyze key components in heterogeneous FL, including data, model, task, communication, and device heterogeneity. For each type of heterogeneity, we discuss in detail the algorithmic challenges it raises.
  2. 2.
    We categorize existing state-of-the-art approaches into data-level, model-level, and architecture-level categories based on the level of processing, and discuss these classifications in depth.
  3. 3.
    We summarize the current privacy-preserving strategies in heterogeneous FL and explore possible future research directions with a view to promoting a safer and more efficient development of the field.
II Classification and Challenges in Heterogeneous Federated Learning
In this section, we first provide a typical federation learning formulation to illustrate the typical federation learning process through an example. Federated learning is a machine learning setup in which multiple clients collaboratively train a model while protecting data privacy [24]. Each participant trains the model on local data and then sends only updates to the model (e.g., weights or gradients) to a central server. The server aggregates these updates and sends the aggregated results back to the participants to update their local models in the next iteration. In this way, federated training produces a global model that performs well on the entire dataset without sharing any raw data. In a classic FL framework [2], assuming that there are $K$ clients participating in the training, for $k$-th client $C_{k}$ has a private dataset $D_{k}=\left\{(x_{i}^{k},y_{i}^{k})\right\}_{i=1}^{N_{k}}$ that satisfies $|x^{k}|=N_{k}$ and $N=\sum_{k=1}^{K}N_{k}$. The local distribution of the data of client $k$ is denoted as $P_{k}(x,y)$, where $x$ and $y$ denote the features of the data samples and the labels. Specifically, FL can be formally defined as the following optimization problem:
$\min_{\theta}F(\theta)=\sum_{k=1}^{K}p_{k}F_{k}(\theta)$ (1)
where $\theta$ is the global model parameter. $F_{k}(\theta)$ is the local loss function on client $k$, which reflects the performance of the model on that client. $p_{k}$ is the client weight, the value of which depends on the amount of data on the client and fulfills $p_{k}=\frac{N_{k}}{N}$ and $\sum_{k=1}^{K}p_{k}=1$. Therefore, the training process of FL can be represented as the following steps:
  • Initialization: The server initializes the global model parameters $\theta$ and sends them to all participating clients.
  • Local training: Each client $k$ trains its own model using its local data and the received global model parameters $\theta$, and computes updates $\Delta\theta_{k}$ to the model parameters.
  • Parameter aggregation: All clients send their model parameters $\theta_{k}$ or gradient updates $\Delta\theta_{k}$ back to the central server. The server aggregates these updates based on the weights $p_{k}$ of each client to update the global model.
    $\mathrm{Model\ parameters:}\theta\leftarrow\sum_{k=1}^{N}p_{k}\theta_{k},$ (2)
    $\mathrm{\textbf{or}\ Model\ gradient:}\theta\leftarrow\theta+\sum_{k=1}^{N}p_{k}\Delta\theta_{k}.$
  • Model broadcast: The central server sends the updated global model $\theta$ to all clients.
  • Model iteration: Repeat steps 2 through 4 until the stop condition is satisfied.
Although FL has been widely used in various domains [1], real-world FL usually faces heterogeneity in data distribution, model architecture, task requirements, network environments, and hardware devices among participant clients [16]. These heterogeneities pose a number of challenges, including how to effectively deal with differences in data distribution across clients, how to adapt to the computational and storage capabilities of different clients while maintaining model performance, how to balance knowledge sharing with task specificity in multitasking scenarios, how to optimize communication strategies to cope with varying bandwidth and latency conditions, and how to adapt to the energy and resource constraints of different devices. We next analyze the challenges posed by these heterogeneities, the resolution of which is essential to achieve efficient, scalable, and robust FL systems.
II-A Data Heterogeneity
Data heterogeneity refers to the inconsistency of data distribution among clients in FL, which does not obey the same sampling, i.e., the data are not independently and identically distributed (Non-IID) [7, 23]. To explore data heterogeneity, we categorize data heterogeneity from the perspective of client data distribution. Specifically, we distinguish different categories of Non-IID data by five different skew patterns: distribution skew, label skew, feature skew, quality skew, and quantity skew, as shown in Fig. 1. Numerous studies[25, 26] have found that the client’s local optimization objective is inconsistent with the global optimization objective due to differences in the client’s local data distribution. As a result, data heterogeneity may cause the local model to converge in different directions to reach the local optimum rather than the global optimum, which reduces the FL performance and leads to poor global convergence of the model, and these problems deteriorate the performance of the global model for each participant, and may prevent participants from engaging in FL [27].
Figure 1: Illustration of five different skew patterns in data heterogeneity.
II-A1 Distribution Skew
In FL, data distribution skew refers to the skew in the distribution of data held by different clients, i.e., the distribution of features or joint distribution of data from different clients is inconsistent with the overall data distribution [28]. In this case, the data no longer obeys the assumption of independent identically distributed (IID). As shown in Fig. 1a, there is $P_{i}(x,y)\neq P_{j}(x,y)$ due to the difference in the data distribution of clients $i$ and $j$.
A typical example is data obeying a Dirichlet distribution [29]. The Dirichlet distribution is a multivariate probability distribution that is often used to describe the distribution of the proportions of each category in a multicategorization problem. In the context of FL, the Dirichlet distribution can be used to model differences in the distribution of data labels on different clients. Where the parameter $\beta$ determines the degree of distributional difference of the data. As shown in the scatterplot in Fig. 1a, which illustrates the visualization of the difference in the distribution of data across 20 clients for different values of $\beta$.
II-A2 Label Skew
Data label skew is a common Non-IID case [30], which refers to the skew in the distribution of data labels held by different clients. This label distribution skew means that even if the feature distributions are shared, the label distributions of different clients may be different [31], i.e., for clients $i$ and $j$, their label distributions $P_{i}(y)\neq P_{j}(y)$. However, for each label, the feature distributions of the data samples are similar regardless of which client they belong to, i.e., $P_{i}(x|y)=P_{j}(x|y)$. For example, in the data label skewing scenario shown in Fig. 1b, different users contain data from the set of handwritten digits, but each client may mainly contain a certain number of digits. The scatterplot shows a visualization of label skew for 20 clients. This data label skew can lead to a large difference in the gradient of the model uploaded back from different clients during federated training, thus affecting the convergence and generalization performance of the model.
II-A3 Feature Skew
Data feature skew is a kind of data heterogeneity in FL, which refers to the inconsistent distribution of data features on different clients [32, 33], e.g.,$P_{i}(x)\neq P_{j}(x)$. In this scenario, even though the labels are consistent $P_{i}(y)=P_{j}(y)$, the distributions remain different $P_{i}(x|y)\neq P_{j}(x|y)$. As in Fig. 1c, clients $i$, $j$, and $k$ all have the label of dog, but they are photos, cartoons, and art paintings of dogs, respectively. This data feature skew is especially common in Vertical Federated Learning (VFL) scenarios [34, 35]. VFL involves multiple clients that have the same set of sample IDs, but each client holds a different set of features. In other words, each client only knows a portion of the features of the samples. For example, consider two financial institutions, A and B, that collaborate to build a credit scoring model. Institution A holds the basic information and transaction history of its clients, while Institution B holds the loan history and repayment status of its clients. Although these two institutions share the same customer base, they hold different sets of data features. In this longitudinal FL scenario, data feature skew manifest itself in the form of different feature spaces held by agencies A and B.
II-A4 Quality Skew
Data quality skew refers to the difference in data quality on different clients [36, 37]. This can be caused by a variety of factors as shown in Fig. 1d, such as noise in the data acquisition process, the data processing flow of different clients, or the accuracy of data labeling. Data quality skew can have a significant impact on the performance of FL models [38]. In this scenario, some clients may have high-quality data that has been accurately measured and carefully labeled, while others may contain a lot of noise or inaccurate labels. High-quality data can provide more accurate information that helps the model learn effective feature representations and prediction rules. On the contrary, low-quality data may introduce misleading information that causes the model to learn incorrect feature relationships, thus reducing the model’s generalization ability and prediction accuracy [39].
II-A5 Quantity Skew
Data quantity skew refers to a significant difference in the amount of data on different clients [40, 41]. For example, some clients may have a large number of data samples while others have only a small amount of data as shown in Fig. 1e. This quantity skew affects the training process of a FL model because the model may tend to optimize the performance on clients with a larger amount of data while ignoring clients with a smaller amount of data. The impact of data quantity skew on the model is mainly in two aspects [42]: on the one hand, clients with larger data quantities may have a larger impact on the global model update, which results in the model performing better on these clients and worse on clients with smaller data quantities. On the other hand, the imbalance in the amount of data may result in the model not being able to fully learn the features of the data on all clients, thus reducing the generalization ability of the model.
II-B Model Heterogeneity
In the model heterogeneity scenario of FL, each client involved in learning may employ different model structures and optimizers to adapt to its own data characteristics and computational resources [8, 43], as shown in Fig. 2. This model heterogeneity increases the complexity of FL systems, but also provides flexibility to adapt to diverse device and task requirements. We will next address model structure heterogeneity and model optimizer heterogeneity, respectively.
Figure 2: Illustration of federated model heterogeneity.
II-B1 Model Structure Heterogeneity
Model structure heterogeneity refers to the fact that different clients use different model structures to adapt to their respective data characteristics and computational resources [44]. This heterogeneity mainly stems from the diversity of devices and the diversity of task requirements in the real world. For example, in a FL system, some clients may be servers equipped with high-performance GPUs capable of handling complex deep learning models such as deep convolutional neural networks (CNNs) or transformer models for image recognition or natural language processing tasks. While other clients may be resource-constrained edge devices, such as smartphones or IoT sensors, which may only be able to run lightweight models, such as shallow neural networks or decision trees, to reduce computational and storage overhead.
Moreover, even for the same type of task, different clients may choose different model structures based on the characteristics of their local data [45]. For example, in the task of performing handwritten digit recognition, some clients may use a standard multilayer perceptron (MLP) model, while others may employ a convolutional neural network model with spatial feature extraction capabilities. This structural difference results in different dimensions and properties of the model parameters, making it difficult to directly apply traditional FL algorithms, such as the Federated Averaging Algorithm (FedAvg), which typically assume that all clients use the same model structure.
II-B2 Model Optimizer Heterogeneity
Optimizer heterogeneity refers to the fact that different clients use different optimization algorithms or parameters to train the model [46, 47]. Due to the different computational power and memory constraints of different devices, some clients may use optimizers with smaller memory footprints and computational costs, such as Stochastic Gradient Descent (SGD), while others may use more sophisticated optimizers, such as Adam or RMSprop, to achieve faster convergence. Moreover, even with the same optimizer, different clients may employ different learning rates or other hyperparameter settings to suit their local data distribution and training conditions. This optimizer heterogeneity increases the complexity of FL, as coordination between different optimization strategies is required to ensure effective global model updating and convergence.
II-C Task Heterogeneity
In FL, task heterogeneity refers to the fact that clients involved in the learning process may face different task requirements or objective functions [9]. This heterogeneity may stem from the environment in which the clients are located, differences in user behavior, or specific application scenario requirements, resulting in different objective functions or learning tasks that need to be optimized for each client.
For example [48], in a FL application in the medical domain, different hospitals may focus on different types of disease diagnosis, so their learning tasks and objective functions may be different, e.g., Fig. 3a demonstrates that clients have homogeneous models but the distributions have labels for different medicines. Fig. 3b demonstrates when clients have similar datasets such as face dataset, and individual clients have heterogeneous classification labels. Fig. 3c Task types are heterogeneous, including classification and regression tasks.
Task heterogeneity introduces additional complexity to FL because traditional FL frameworks typically assume that all devices jointly optimize a global objective function [49]. To cope with task heterogeneity, FL systems need to employ more flexible algorithms and policies that allow different devices to personalize model training based on their own task requirements. In addition, task heterogeneity also requires that FL systems are able to effectively coordinate the learning processes of different devices to ensure a balance between the generalization capabilities of the global model and the individual needs of each client. Such coordination may require appropriate adaptation of the model updating and parameter aggregation processes to suit the characteristics and needs of different tasks.
Figure 3: Illustration of federated task heterogeneity.
II-D Communication heterogeneity
In FL, communication heterogeneity refers to the differences in communication capabilities of the devices involved in the learning process [10, 50]. Such differences are mainly in terms of network bandwidth, latency and connection stability, which directly affect the transmission efficiency of model parameters and the synchronization of the whole learning process. For example, devices are usually deployed in different network environments with different network connection settings (3G, 4G, 5G, Wi-Fi) [51, 52]. Devices with high bandwidth and low latency can quickly upload and download model parameters, thus completing a round of model updates in a short period of time. On the contrary, devices with limited bandwidth and high latency may face longer waiting time when transmitting model parameters, resulting in the training progress of the whole system being affected. The heterogeneity of communication increases the communication cost and complexity to some extent [53]. In addition, differences in connection stability can also have an impact on the efficiency of FL. Unstable connections may result in lost or delayed model parameter updates, which in turn affects the convergence speed and accuracy of the model.
II-E Device heterogeneity
In the application scenario of FL, device heterogeneity is a crucial issue that cannot be neglected [54, 55]. Due to the wide variety of client devices participating in FL, there are significant differences in computing power, storage resources, energy supply, etc., which leads to device heterogeneity [11]. This heterogeneity leads to inconsistent efficiency and capability of different devices in processing data and training models, which may affect the performance and efficiency of the whole FL system. For example, high-performance devices such as servers or high-end smartphones are able to process complex models and large data volumes quickly due to their powerful processors and ample storage space, thus taking on more computational burden during FL and speeding up the entire learning process [56]. On the contrary, low-performance devices, such as low-end smartphones, are less capable of model training and data processing due to the limitations of computational resources and storage space, and may also slow down the progress of the entire FL system, which may become the bottleneck of system efficiency. This is because the system usually needs to wait for all devices to complete their local updates before global model aggregation can take place. In addition, devices with weaker computational power may face greater pressure on energy consumption when processing complex models, which is especially critical in energy-constrained environments.
Therefore, FL algorithms and system design need to take into account the heterogeneity of computational capabilities among devices, and ensure efficient operation and fairness of the system through strategies such as reasonable task allocation, model simplification, and resource optimization to ensure that all devices can effectively participate in the learning process, while maintaining the coordination and fairness of the overall system.
III Advanced Methods Discussion
III-A Data-Level Methods
Doing operations at the data level is often seen as an important means to address data heterogeneity. In this section, we categorize the data-level approaches into three categories: data cleansing, data augmentation, and representation learning, and present their representative approaches respectively.
III-A1 Data Cleaning
Data cleaning in FL can help improve the training effect and final performance of the model. Data cleaning mainly refers to correcting errors and inconsistencies in datasets, including dealing with issues such as missing values, outliers, format errors, and duplicate data.
Li et al. [57] presents an optimization problem i.e., selecting a high-quality set of training samples for a FL task under a given budget in a privacy-preserving manner without knowledge of the participants’ local data and training process. Ma et al. [58] proposes a federated data cleaning protocol, called FedClean, for edge intelligence scenarios that aims to achieve data cleaning without compromising data privacy. FedClean Different edge nodes first generate a Boolean share of their data and distribute it to two clean servers. These two servers then run the FedClean protocol to compute and sort the data entry scores. Faced with the problem of noisy data labels, Chen et al. [59] proposes a robust FL by considering two networks with different data distributions. The core idea is that the edge devices learn to assign local weights of the loss function in the noisy labeled dataset and cooperate with the central server to update the global weights as a way to help edge nodes that do not have a clean dataset to reweight their noisy loss functions. Similarly, Duan et al. et al. [60] proposes a new data filtering approach to deal with label noise in joint learning, called FED-DR-Filter, which focuses on identifying clean data by exploiting the correlation of global data representations. Tsouvalas et al. [61] proposes FedLN framework enough to deal with labeling noise in different FL training phases, i.e., FL initialization, on-device model training, and server model aggregation, which is able to adapt to the different computational power of the devices in the FL system. Li et al. [62] proposed FedDiv to address the challenge of FL with noisy labels, and they designed a global noise filter with a joint noise filter for efficiently identifying samples with noisy labels on each client, which improves the stability of the local training process. Xu et al. [63] proposed a collaborative data filtering method Safe to deal with poisoning attacks, using the alternating direction method of multipliers to detect the attacked devices whose training process will be interrupted, and using the K-mean clustering algorithm to cluster the trusted data to find out and filter out the poisoned data from the attacked devices.
III-A2 Data Augmentation
Data augmentation plays a key role in FL, especially when dealing with Non-IID data distributed across multiple devices. Data augmentation not only increases the amount of data, but also increases the diversity of the data.
Hao et al. [64] proposes a FL system and derives 2 variants that can use zero-sample data augmentation on underpresented data to mitigate data heterogeneity. Chen et al. [65] proposes FRAug, a federated representation augmentation method, to address the data heterogeneity challenge. FRAug optimizes a shared embedding generator to obtain client consensus, and then locally transforms the output synthetic embeddings to client-specific embeddings in order to increase the training space for each client. Lewy et al. [66] proposed StatMix, an enhancement method using image statistics to improve the results of FL scenes. Xin et al. [67] proposes FL-GAN, a generative adversarial network model for differential privacy based on joint learning. by combining the Lipschitz limit with differential privacy sensitivity, the model can generate high-quality synthetic data without sacrificing the privacy of the training data. Zhang et al. [68] proposes a distributed data augmentation method that combines generative adversarial networksand federated analysis to enhance the generalization ability of trained FDG models, called FA-FDG. Generating more overlapping data by learning the features of limited overlapping data and many locally existing non-overlapping data, Zhang et al. [69] proposed a data augmentation method FedDA based on generative adversarial networks to increase the amount of training data. Similarly, Xiao et al. [70] proposes FeCGAN, a distributed generative adversarial network for multiple participants with insufficient data overlap.
III-A3 Representation Learning
Representation learning aims to learn useful data representations (or features) from dispersed data sources to help improve the predictive performance and generalization of models.
To address the problem of data heterogeneity, Liu et al. [71] introduces federated transfer learning (FTL). FTL does not compromise user privacy and can transfer knowledge across domains in a data federation. This enables the target domain to utilize the rich labels of the source domain to build flexible and efficient models. Li et al. [72] proposes FedMask, a FL framework for efficient communication and computation, where each device can learn an efficient personalized structured sparse model. Zhang et al. [73] proposes for Federated Unsupervised Representation Learning (FURL) to learn generic representation models without supervision while preserving data privacy. FURL solves the problem of inconsistent and misaligned representation space through a dictionary module and an alignment module. Similarly, Yan et al. [74] uses unsupervised representation to pre-train deep neural networks in a federated environment using unlabeled data, which allows for the extraction of discriminative features using pre-trained networks. Wu et al. [75] proposes a communication-efficient FL algorithm, FedOnce, which is based on the privacy-preserving technique of momentary accounting and employs an unsupervised learning representation that enables better results to be achieved with only one-time communication between the parties. Li et al. [76] proposed MOON to correct for local training of individuals using similarities between model representations. Jang et al. [77] proposes a personalized federated learning approach that applies local feature representation learning to stabilize the decision boundary and improve the local feature extraction capability of the client. Tan et al. [78] proposes FedProto, which only passes the prototype to the client without passing model parameters and gradients.
III-B Model-Level Methods
In this section, we discuss some advanced model-level FL methods for solving various heterogeneity types. Model-level methods means that these methods improve FL in terms of model parameters, model output, etc. Specifically, we classify these methods based on their adopted techniques, including parameter decoupling, model compression, knowledge distillation and federated optimization.
III-B1 Parameter Decoupling
Parameter Decoupling based methods regard the whole model parameters as the composition of components, which play different roles in the process of training or inferring. Depending on the goal of parameter decoupling, different methods usually have different granularities. Some methods only regard the whole model as the composition of a backbone model and a head model, while some others decouple the model in a layer-wise or channel-wise, even coordinate-wise manner. Arivazhagan et al. [79] propose FedPer, which split global model into base layers and personalized layers. Base layers are layers closer to the input side while the personalized layers are the remaining layers, and each client can customize the own architecture of personalized layers while sharing the architecture of base layers. In each round, only the parameters of base layers are uploaded to the server and then aggregated. Based on the belief that data from all parties involved in FL often share global feature representations while reflect distribution and label skew in the labels, Collins et al. [80] propose FedRep, which separate model into representation part and head part. The clients train a global representation in a federated manner and then train their own head based on the learned representation. Luo et al. [81] empirically verify that distribution and label skew impact more on deeper layers, and then propose CCVR. CCVR utilizes regularization and calibration in local training to reduce weight bias, and trains a GMM to generate virtual representations for further classifier training. To solve the problem of long-tailed data in FL, Shang et al. [82] propose CReFF, where clients upload real feature gradients, and then the server use them to optimize federated feature gradients for re-training classifier. Comparing to above works that enable shared global representations of features, Liang et al. [83] propose LG-FedAvg, where clients share a common architecture of classifier, and each client learns compressed local representation for training a classifier suitable for local representations of all clients. Learning such local representations on the clients can also help alleviating task heterogeneity, encouraging the clients to learn and extract representations suitable for global classifiers. Xu et al. [84] propose FedPAC, which encourage each client to learn global task variant information by regularizing local representations to approach the global feature centroid. In order to train personalized client models under feature skew, distribution skew and other scenarios, Shen et al. [85] propose CD2-pFed, which implement channel-wised personalization by channel-decoupling, and utilize personalized-weights to dynamically control the degree of personalization in each layer. Besides, CD2-pFed also utilizes cyclic-distillation to mutually distill local and global model representations, bridging the semantic gap between the decoupled channels. Recently, Chen et al. [86] suggest that biased projection heads may cause unreliability on models when faced with data heterogeneity. Therefore, they propose APH method, where clients can get multiple initialized parameters from existing projection heads, and then fine-tune and average them, thus eliminating the bias.
In order to reduce the memory cost of clients, Pillutla et al. [87] propose to divide the model parameters into two parts: shared parameters and personal parameters. Therefore, they propose FedSim and FedAlt, representing training both part of parameters in a simultaneous manner or an alternate manner, respectively. Yi et al. [45] propose personalized FL method FedGH, where clients share a generalized head. The clients upload local average representation, which is utilized for individually training of the generalized head on the server side. Recently, to achieve model personalization when it is model heterogeneity, Wang et al. [88] propose pFedHR. In each round, the server performs layer-wise decomposition on client’s models, and then groups the layers. After that, the server uses the grouped layers to reassembly candidate models and fine-tune them with a public dataset.
To alleviate the impact of device heterogeneity and communication heterogeneity, Diao et al. [51] propose HeteroFL, where the server determines the complexity level of each client based on its computation and communication capabilities, and assigns sub-models of different sizes to clients according to their complexity level. Recently, Yi et al. [89] point out that even pruned models can also reveal the real architecture of the global model. Therefore, they propose a novel framework FedP3, where clients only upload selected model parameters back to the server for privacy protection. Pfeiffer et al. [90] empirically observe that the techniques based on sub-models suffer from performance degradation because the co-adaptation between parameters is broken. Thus, they propose Successive Layer Training (SLT). All clients can train the same parameters, but they begin at early layers with a down-scale head. After early layers are well trained, then SLT squeeze the parameters of early layers and expand the head layers for further training.
III-B2 Model Compression
Model compression refers to compressing the size of a model and is commonly used to solve model, device and communication heterogeneity issues. The commonly used compression methods in FL include sparsification, quantization and factorization. Sparsification usually means identify and retain the meaningful enough parameters, while setting other parameters to some constants. In this way, the parameter matrices are transformed to be sparse, and only those meaningful enough parameters are transmitted. Different from sparsification, quantization keeps the parameters dense and focuses on reducing the bit-width occupation of each parameter. A representative quantization method is SignSGD [91], which only retains the sign bit of each parameter. Factorization decompose a large matrix into the product of multiple low-rank matrices or vectors, and thus compress the size of the whole model with limited precision degradation. To compress both upstream and downstream communications, Sattler et al. [92] point out that sparsification is more robust to Non-IID data and thus propose Sparse Ternary Compression (STC), which further quantizes the remaining top-$k$ elements to the mean population magnitude. Similar to the goal of [92], Amiri et al. [93] propose Lossy Federated Learning (LFL). The server broadcast the quantized version the global model parameters to the clients, and the uploading parameters are also quantized. Shah et al. [94] propose to compress both downstream and upstream communications, but it goes further. To recover the loss of sparsity after aggregation, each client can perform sparse reconstruction for the averaged model, and it can also directly begin local training with the received sparse model, which can help in inducing sparsity in the client model. Recently, Tang et al. [95] point out that SignSGD is vulnerable to Non-IID data, and current corresponding methods cannot support multiple local SGD updates like FedAvg. Thus, they propose to inject some random noise before the sign operation to control the bias brought by it, and further propose the first sign-based FedAvg algorithm $z$-SignFedAvg.
To combat the challenge of model and device heterogeneity, Yao et al. [96] propose FedHM, where the server compresses the large global model via low-rank factorization for each client to satisfy its resource availability, while preserving the model’s performance. Moreover, the compressed model can be transformed back to full-rank model for aggregation. Jeong et al. [97] propose Factorized-FL, which aims to factorize model parameters into basis vectors. Such factorized parameters space can make aggregation operation more robust to different labels and data, and separate client-specific and client-general knowledge. Recently, to perform personalized FL under scenario with data heterogeneity and device heterogeneity, Chen et al. [98] propose a novel personalized FL framework Fed-QSSL, where clients can perform low-bitwidth quantization training locally to satisfy its device constraint, and tackle data heterogeneity with distributed self-supervised learning.
III-B3 Knowledge Distillation
Knowledge distillation [99] is originally proposed for transferring knowledge from a larger model to a smaller model, and thus reducing the number of parameters. Generally speaking, This process requires the output of the smaller model to approach the larger one. Therefore, knowledge distillation can also be used for knowledge transferring between participants in FL. To alleviate the impact of feature skew and model heterogeneity, Huang et al. [100] propose FCCL, where each client constructs a cross-correlation matrix according to other clients’ soft labels on public dataset. Besides, FCCL also utilize dual-domain knowledge distillation to alleviate catastrophic forgetting caused by domain shift. Zhu [101] et al. propose FedGEN, where the server trains a light-weighted representation generator, and broadcast it to clients. While local training, the clients utilize the generated representations for knowledge distillation, hence get rid of reliance on public datasets. For scenarios with model heterogeneity and quality skew, Fang et al. [38] proposes RHFL, where clients combine cross entropy and reverse cross entropy to alleviate the impact of label noise, and perform knowledge distillation on a public dataset to reduce the knowledge gap between clients. Besides, RHFL also introduces client confidence re-weighting to reduce the negative impact caused by other clients’ noise. Recently, to combat severe label noise because of highly contaminated clients, Lu et al. [23] propose FedNed, which identifies noisy clients and requires them to train models with noisy labels and pseudo-labels simultaneously. Then the models trained with noisy labels act as bad teachers for negative distillation, while the ones trained with pseudo-labels are used for aggregation. To learn personalized models under data heterogeneity, Chen et al. [102] propose a method, which captures similarity between generic model and personalized models by spectral distillation. Moreover, it also establishes a co-distillation framework to bridge the knowledge of the generic model and personalized models.
III-B4 Federated Optimization
Some methods introduce other techniques to the process of federated optimization, which can help the models to more adaptable to the local distributions or the global distribution. We mainly discuss three federated optimization techniques applies in model-level approaches: Multi-task learning, meta-learning and regularizing.
Federated multi-task learning. The core idea of multi-task learning is to train a model simultaneously to perform multiple related tasks, thereby improving the model’s generalization ability. By regarding each client’s local training as a separate task, multi-task learning techniques can help clients to learn personalized model with high generalization ability. Smith et al. [9] extend MTL method CoCoA from traditional distributed learning to federated environment, and propose MOCHA, which considers each client’s local training as a task and train a personalized model for each client. Corinzia et al. [103] propose to construct a Bayesian network between the server and the clients, and utilize variational inference for its optimization. Comparing to MOCHA, it requires no additional regularization term in local training objective and thus improve the efficiency of local training and ensuring network-agnostic FL. Besides, it also extends FMTL to scenarios with non-convex objectives. To learn a personalized model for devices that later join FL, Li et al. [104] propose OFMTL, which helps the new devices to learn with the assistance of weight matrix and precision matrix constructed by old devices, requiring no presence of earlier devices. Dinh et al. [105] propose FedU, which introduce Laplacian regularization to determine the correlation between clients in FMTL. Besides, They also propose a decentralized version dFedU. Marfoq et al. [106] consider the data distribution of each client as a mixture of several latent distributions. Therefore, they propose FedEM, where clients can collaboratively train several generic component models, and each client only needs to learn how to mix these components to build a personalized model.
Meta-learning based FL. Meta-learning, which is also known as ”learning to learn”, is usually used to design and train algorithms to quickly adapt to new tasks, typically when it is very little data. The core goal of meta-learning is to improve the flexibility and adaptability of learning algorithms, enabling them to learn from previous experiences and better handle new and unknown tasks. Chen et al. [107] combine FL with meta-learning and then propose FedMeta, where all clients collaboratively train a initialization model. In FedMeta, each client upload its test loss on its query set which is used to update the initialization model. Each client can learn a personalized model from the initialization model by just a few local training, thus FedMeta can also reduce the training burden of the clients. Recently, Jeon et al. [108] propose a Bayesian meta-learning based approach called MetaVD, where the server establishes a shared hypernetwork for predicting client-specific dropout rates, which are utilized for few-shot local adaptation on clients. Leo et al. [109] propose FedL2P, which introduces meta-nets to inductively map each client’s Batch Normalization statistics to client-specific personalization hyperparameters. Scott et al. [110] propose PeFLL. The server is equipped with a hypernetwork, which takes the descriptor of each client as input and personalized model parameters as output. The descriptor of each client is obtained with an embedding network, taking some data points as input. Thereby PeFLL overcomes the problems of high latency and client-side computation cost.
Regularization based FL. Regularization usually means adding a regularization term in the local training objectives and thus achieving some specific goals, such as tackling data heterogeneity. Li et al. [43] propose FedProx, which adds a proximal term in local training objective to reduce the distance between the local model and the global model, and thus handles data heterogeneity. Besides, They also propose the concept of in-exact solution, enabling the clients to upload incompletely solved parameters and thus alleviate the problem of device heterogeneity. Kim et al. [111] propose FedMLB, which separate models into several blocks, and each client form a main pathway and several hybrid pathways with the blocks. Regularization terms based on knowledge distillation are formed with the output of the main pathway and the hybrid pathways. Shoham et al. [112] point out that due to data heterogeneity, the problem of forgetting in Sequential Lifelong Learning may also occur in FL. Therefore, they propose FedCurv, which introduces Elastic Weight Consolidation (EWC) to regularize model parameters during local training, and thus avoid forgetting the learned information from other clients. Similarly, Yao et al. [113] propose FedCL, which utilize a proxy dataset to construct importance weight matrix of all local models to perform EWC regularization, and thus alleviate weight divergence. Dinh et al. [114] propose pFedME, which introduce Moreau Envelope for regularization. It adds an L2 regularization term to the local objective to transform the entire FL into a two-level objective. The inner optimization obtains optimal personalized models while the outer one obtains the global model. Karimireddy et al. [115] propose SCAFFOLD, utilizing control variables that reflect the update direction of the global model and the local models respectively, and then correct update direction of local model with a regularization term formed by their differences. Hanzely et al. [116] propose a new formulation of FL, which no longer learn a single global model, but a mixture of the global model and the pure personalized models. To perform such FL formulation, this works proposes a novel randomized gradient-based method L2GD, for minimizing the loss function composed of average loss and penalty term. Li et al. [117] propose a personalized FL method Ditto, where each client perform local training on the global model and then on its personalized model, and applies regularization to the personalized model’s training to bring it closer to the global model. Besides, They also theoretically and empirically prove that Ditto has good robustness and fairness guarantee. Recently, Zhou et al. [118] propose FedLNL, which designs a novel method based on Bayesian Inference for updating local Noise Transition Matrix (NTM) and thus avoid too large gradients, and then apply a diversity product regularizer in local training to utilize this method. Zhi et al. [119] suggest that each client can guide its local training referring to other clients’ parameters, and thus propose a regularizer based on a relation cube to conduct layer-wise parameters coaching. Li et al. [120] empirically observe that current FL algorithms are highly susceptible to the degree of weight decay, and propose FedNAR. FedNAR is an algorithmic plug-in which can help controlling the magnitude of weight decay.
III-C Architecture-Level Methods
To address the federated heterogeneity challenge, making adjustments in the training structure is an emerging improvement. In this section, we discuss some heterogeneous methods that make adjustments on the FL architecture. Specifically, we classify these structure-level approaches into client selection, clustered federated learning, hierarchical federated learning, and decentralized federated learning.
III-C1 Client selection
Client selection refers to the process of selecting a subset of available clients to participate in model updates during the training process. This selection can directly impact the efficiency and effectiveness of the training. Effective client selection can address issues of data heterogeneity and constraints related to communication and computation [121, 122].
To alleviate the pressure of heterogeneity, Nishio et al. [53] proposed that FedCS uses a greedy algorithm to solve the client selection problem with resource constraints, which allows the server to aggregate as many client updates as possible and accelerate the performance improvement of ML models. Yoshida et al. [123] adopted a heuristic algorithm and proposed a hybrid learning mechanism called Hybrid-FL, trying to select the best set of clients to use their own data to train the model. Wang et al. [7] introduced an experience-driven control framework, Favor, which intelligently selects client devices to participate in each round of FL to offset the bias introduced by non-IID data and accelerate the convergence speed. Xu et al. [124] proposes a stochastic optimization problem that jointly connects client selection and bandwidth allocation under long-term client constraints. lai2021oort et al. [125] proposed Oort to improve the performance of joint training and testing by guiding participant selection. Oort uses existing data to improve model accuracy and speed up training. Wu et al. [126] proposed a probabilistic node selection framework, FedPNS, which dynamically changes the probability of each node being selected based on the output of the optimal aggregation. Cho et al. [127] presents a convergence analysis of FL of biased customer selection via efficient selection aggregation and quantifies how bias affects the convergence speed. Li et al. [128] proposed PyramidFL to speed up FL training while achieving higher final model performance. At the heart of PyramidFL is fine-grained client selection, not only focusing on the differences in client selection between those selected and unselected participants, but also fully exploiting data and system heterogeneity within selected clients. Ribero et al. [129] proposes a federated averaging algorithm assisted by an adaptive sampling technique, F3AST, an unbiased algorithm that dynamically learns an availability-dependent client selection strategy that asymptotically minimizes client sampling variance to global model convergence The impact improves the performance of joint learning.
III-C2 Clustered Federated Learning
Clustering FL is used to deal with situations where the data distribution is highly heterogeneous. This approach trains the model in a more homogeneous group by clustering clients based on the similarity of their data.
CFL (Cluster Federated Learning) [130] is a method of recursively clustering clients by exploiting the cosine similarity between client gradient updates during each round of FL training. This clustering allows for a more targeted and efficient distribution of training tasks among clients with similar characteristics. IFCA (Intra-cluster Federated Average) [131] is another clustered FL method that utilizes the loss of models within different clusters on the client’s local dataset to cluster clients. In each round, IFCA randomly selects a portion of clients within each cluster for training and updates the intra-cluster models by aggregating the local models within the cluster. PACFL (Personalized and Clustered Federated Learning) [132] clusters clients by uploading data information from each client to the server and taking measurements of the principal angles between client data subspaces before the FL process begins. During each round of joint learning, PACFL updates each intra-cluster model with a weighted aggregation of the local models of the intra-cluster clients, enabling personalized learning while still taking advantage of clustering. FedSoft [133] is a cluster FL framework designed to train both local personalized models and high-quality cluster models. In order to limit the workload of the clients, FedSoft employs proximal updating, which allows only a subset of clients to complete an optimization task in each round of communication. ACFL (Auction-based Cluster Federated Learning) [134] introduces a mean-shift clustering algorithm that intelligently groups clients based on their local data distributions. This clustering approach allows FL systems to identify groups of clients with similar data characteristics, enabling more targeted and efficient model training.
III-C3 Hierarchical Federation Learning
Hierarchical Federated Learning is a more organizationally complex FL architecture that employs a multilevel federation model to manage and train clients distributed at different levels. This model is usually suitable for scenarios with a multi-layered network structure. It is worth noting that hierarchical FL focuses on how to organize and optimize the learning process in networks with a well-defined hierarchical structure, e.g., federated edge learning. In contrast, clustered FL focuses more on grouping clients based on data similarity without the need for an explicit hierarchical structure.
Liu et al. [135] propose a client-edge cloud hierarchical joint learning system, HierFAVG, an algorithm that allows multiple edge servers to perform partial model aggregation, which allows for better communication and utility tradeoffs. Abad et al. [10] collaboratively learn global models by sharing local updates of model parameters instead of their dataset, and in this way use a hierarchical joint learning scheme to significantly reduce communication latency without sacrificing accuracy. Briggs et al. [136] separated clusters of clients by introducing a hierarchical clustering step that uses the similarity of the client’s local updates to the global joint model. Xu et al. [137] consider a hierarchical FL system and propose a joint problem of edge aggregation interval control and resource allocation to minimize a weighted sum of training loss and training delay. Lim et al. [138] use an evolutionary game to model workers’ edge association strategies and propose a hierarchical game framework to study the dynamics of edge association and resource allocation in self-organizing HFL networks. Abdellatif et al. [52] investigated the potential of hierarchical FL in IoT heterogeneous systems and proposed a user assignment and resource allocation method on a hierarchical FL architecture for IoT heterogeneous systems. Zhou et al. [139] proposed a unique clustering-based approach for participant selection using social context data. By creating different groups of edge participants and performing group-specific joint learning, the models of various edge groups are further aggregated to enhance the robustness of the global model. Ma et al. [140] proposed FedUC, which clusters workers through a hierarchical aggregation structure to solve the edge heterogeneity problem and reduce the communication consumption of FL. Deng et al. [141] proposed a communication cost minimization problem in the HFL framework by making decisions on edge enhancer selection and node edge association in order to achieve the target accuracy of the total communication cost required for model learning.
III-C4 Decentralized Federated Learning
Decentralized federated learning (DFL) is an approach that does not rely on a central server to manage model training. Compared with traditional centralized federated learning, decentralized methods provide higher privacy protection.
Li et al. [142] developed a decentralized FL framework BFLC based on blockchain with committee consensus. It uses blockchain for global model storage and local model update exchange [143], and reduces malicious client attacks through committee verification composed of honest clients. Lim et al. [144] proposed a deep learning-based auction mechanism to derive the valuation of the services of each cluster leader. ye et al. [145] proposed a robust decentralized stochastic gradient descent federated method, called soft DSGD, to solve the problem of unreliable transmission process. Pappas et al. [146] designed a fully decentralized FL framework based in part on the Inter Planetary File System (IPFS). By using IPLS and connecting to the corresponding private IPFS network, any party can initiate the training process of a machine learning model, or join an ongoing training process initiated by the other party. Liu et al. [147] proposed a general DFL framework that regularly implements multiple local updates and multiple inter-node communications to strike a balance between communication efficiency and model consistency. Kalra et al. [148] proposed an efficient communication scheme for decentralized FL, called ProxyFL. Each participant in ProxyFL maintains two models, one is a private model and the other is a public shared proxy model. Designed to protect participant privacy. The agent model allows efficient information exchange between participants without the need for a centralized server. HegedHus et al. [149] believe that Gossip learning is a decentralized alternative to FL, so they designed a DFL training algorithm based on the Gossip protocol. Qu et al. [150] proposed a new architecture called DFL for drone networks (DFL-UN), which implements FL in drone networks without the need for a central entity. Beltran et al. [151] introduced a novel platform, Fedstar, which allows users to create federations by customizing parameters. Fedstar aims to train FL models in different federations of physical or virtual devices in a decentralized, semi-decentralized and centralized manner. Zhou et al. [152] proposed a decentralized federated trusted average (DeFTA), as a plug-and-play replacement for FedAvg, brings immediate improvements to security, scalability, and fault tolerance in the FL process. Cheng et al. [153] used a decentralized and trusted aggregation strategy to introduce DeTA, a decentralized FL system architecture, and designed a defense-in-depth model.
IV Privacy Protection in Heterogeneous Federated Learning
In this section, we will focus on the privacy and security issues in heterogeneous FL. As is well recognized, FL frameworks are deployed over the internet. The complex and unpredictable network environment poses potential vulnerabilities for attacks on the FL process. This underscores the necessity of employing protective measures within FL applications. Since the inception of the concept of FL, researchers have published a range of attack methods targeting the exposure of certain stages to the network environment. These methods can launch attacks at various stages, such as dataset preparation [154], model transmission [155], and model aggregation [156], with malicious intentions of privacy theft [157] or interference [158] with the training process. Among them, model inversion attacks [12], which involve eavesdropping on the model transmission channel to intercept the model and attempt to reconstruct the client’s private data from the captured information, pose significant threats to FL frameworks. In some cases, it is even possible to fully reconstruct the private dataset from the model. For attack vectors such as model inversion attacks, model protection is particularly critical. Through years of research, model protection can be categorized into three types [15]: differential privacy, secure multi-party computation, and model tracing, each providing pre-emptive or post-emptive protection for the transmitted models from different perspectives. It is noteworthy that different types of protective methods will have varying impacts on the trade-off [159, 160] between privacy, utility, and efficiency, thus analyzing privacy-preserving strategies is crucial for privacy-utility trade-offs in heterogeneous FL. We show common FL privacy-preserving strategies in Fig. 4.
IV-A Differential Privacy
Differential privacy (DP) [161], initially proposed by Dwork et al., is a privacy definition from the field of database security. Since the introduction of FL, differential privacy has attracted the interest of many researchers due to its provable and quantifiable privacy. Generally speaking, FL frameworks meet the requirements of DP through a noise mechanism [162], which means that almost all DP methods will sacrifice model utility to obtain privacy, with little impact on training efficiency.
Figure 4: Illustration for Protect Methods
IV-A1 DP and Noising Mechanism
Based on relevant derivations, we generally use a noise mechanism to perturb the transmitted model, thereby making the entire FL mechanism satisfy DP. This mechanism is also known as model perturbation. Before introducing the noise mechanism in detail, we first give the formal definition of DP.
Theorem 1 ( $(\varepsilon,\delta)$-DP).
A randomized mechanism $\mathcal{M}:\cal{X}\mapsto\cal{R}$ with domain $\cal{X}$ and range $\cal{R}$ satisfies $(\varepsilon,\delta)$-DP if, for any two adjacent inputs $x,x^{\prime}\in\cal{X}$ and for any subsets of outputs $\cal{S}\subseteq\cal{R}$ it holds that
$\Pr[\mathcal{M}(x)\in\mathcal{S}]\leq\exp(\varepsilon)\cdot\Pr[\mathcal{M}(x^{\prime})\in\mathcal{S}]+\delta$ (3)
where $(\varepsilon,\delta)$ is privacy budget. The smaller the value of $\varepsilon$, the stronger the level of protection provided by differential privacy.
In the context of model perturbation, the model information prepared for transmission undergoes two stages of processing. The first stage is model clipping, which involves setting a threshold $C$. If the norm of the model exceeds this threshold, the model is scaled proportionally until its norm equals the threshold. The second stage involves noising, specifically generating Gaussian noise with a chosen standard deviation $\sigma$ to perturb the model.
[…]
The relationship between the noise amplitude $\sigma$ and the privacy budget $\varepsilon$ can be concluded from relevant DP theories, such as Parallel Composition Theorem, Advanced Composition Theorem [162], and so on. Compared to the theories mentioned above, the Moment Accountant [163] theory provides a smaller upper bound for the privacy budget, meaning that this theory can derive a smaller $\varepsilon$ from the same $\sigma$.
There are numerous federated learning frameworks that apply model perturbation. For instance, DP-FedAvg [164], which was initially proposed, first adopted user-level DP to protect client data. NbAFL[165] considered the noise magnitude required for differential privacy protection of model information in both upload and download links. The work of Seif et al. [166] considered the environment of wireless communication and proposed how to implement differential privacy protection in this environment by utilizing the characteristics of wireless communication. CLDP-SGD [167], while maintaining DP protection, considered communication overhead and is a federated learning architecture that applies DP under communication-limited conditions. In addition, some works like Fed-CDP [168] set the object of protection as the training set data, that is, noise is added during local training.
IV-A2 Variant DP
In traditional DP, the relationship between the magnitude of noise and the privacy budget is intricate and not easily described intuitively. To address this issue, Ilya have slightly relaxed the requirements of DP, resulting in the formulation of Rényi Differential Privacy (RDP) [169].
Theorem 2 ($(\alpha,\rho)$-RDP).
Given that $\alpha\in(1,\infty)$ and $\rho>0$, a randomized mechanism $\mathcal{M}$ with domain $\cal{X}$ and range $\cal{R}$ satisfies $(\alpha,\rho)$-RDP if, for any two adjacent inputs $x,x^{\prime}\in\cal{X}$ the $\alpha$-Rényi divergence satisfies
$D_{\alpha}[\mathcal{M}(x)\parallel\mathcal{M}(x^{\prime})]\leftarrow\frac{1}{\alpha-1}\log\mathbb{E}\left[\left(\frac{\mathcal{M}(x)}{\mathcal{M}(x^{\prime})}\right)^{\alpha}\right]\leq\rho$ (5)
where the expectation is taken over $\mathcal{M}(x^{\prime})$.
RDP offers several advantages for federated learning, which are encapsulated in the following points [169]:
  • It is compatible with the Gaussian mechanism, and it can be proven that when employing a Gaussian noise mechanism, the relationship is given by $\sigma^{2}=\frac{(\Delta f)^{2}\alpha}{2\rho}$, where $\Delta f$ refers to the sensitivity of function $f$ and defined as $\Delta f\leftarrow\max_{x,x^{\prime}\,\text{adjacent}}\left\|f(x)-f(x^{\prime})\right\|_{2}$;
  • RDP is amenable to composition, meaning that the outcomes processed sequentially through $(\alpha,\rho_{1})$-RDP and $(\alpha,\rho_{2})$-RDP mechanisms satisfy the composition property of $(\alpha,\rho_{1}+\rho_{2})$-RDP.
  • RDP can be transformed into the classical DP, facilitating comparative analysis of privacy with other mechanisms.
More recent research in the field of federated learning has replaced classical DP with RDP as the basis for privacy. DP-FedSAM [170] is proposed under the consideration of RDP on how to improve the optimizer to enhance model utility. Fed-SMP [171] considers how to apply RDP’s privacy guarantee to the model after sparsification.
IV-A3 DP with Heterogeneity
In real-world environments, federated learning frameworks with Differential Privacy may encounter several issues due to heterogeneity. We categorize these issues as follows:
  • Different clients have varying requirements for privacy, making it difficult to determine the privacy of the entire system;
  • If a client’s privacy budget is too small, the model information from that client is likely to affect model performance [164];
  • Data heterogeneity itself can lead to a loss in model utility, and DP can make this loss more apparent [172].
In response to these issues, the academic community has proposed some solutions. PFA [173] classifies clients into public and private categories based on their privacy budgets. After the models of the two types of clients are aggregated separately according to the privacy budget, the model of the private client is projected into the space of the public client’s model. This takes into account the different credibility of information with different privacy levels and does not waste data with stronger privacy. In some cases, the network environment where the federated learning algorithm is deployed also has heterogeneity, and FedSeC [174] has proposed improvements from this perspective. There are also works like DP-SCAFFOLD [175], which advance from the perspective of reducing model utility loss.
IV-B Secure Multi-party Computation
Secure Multi-party Computation (SMC) [176] refers to methods that ensure the security of distributed computation in complex network environments, often with security guarantees provided by cryptography. Due to the nature of cryptography [177], the model information processed by SMC is almost undamaged, ensuring that the utility of the model does not decrease. However, these methods often involve complex computational processes [178], leading to significant time and space overhead during the model training process.
IV-B1 Homomorphic Encryption
Homomorphic encryption (HE) [179] is a specialized cryptographic scheme that imposes additional requirements on certain computational properties to be preserved beyond traditional encryption/decryption functionalities. Different types of HE satisfy distinct properties as follows [180]. Assuming that +,+$\cdot$+ , ⋅ are addition and multiplication on plaintext space and $\oplus$,$\odot$ , ⊙ are ones on cipher space, and $\mathsf{E}$ refers to encryption operation.
  • Addictive Homomorphism. $\mathsf{E}(a+b)=\mathsf{E}(a)\oplus\mathsf{E}(b)$ for any plaintexts $a,b$;
  • Multiplicative Homomorphism. $\mathsf{E}(a\cdot b)=\mathsf{E}(a)\odot\mathsf{E}(b)$ for any plaintexts $a,b$;
A cryptographic scheme satisfying only one property is termed partially HE, whereas one that satisfies both properties is named fully HE. When HE is employed in federated learning [181], the model can be encrypted before transmission, preventing even observers or servers from accessing model information. Building upon this foundation, leveraging the properties of homomorphic encryption, servers can perform aggregation operations on the model without acquiring the model information. In the current landscape of common homomorphic encryption schemes, the Paillier cryptosystem [182], as an additive homomorphic encryption scheme, was introduced relatively early and is extensively used in federated learning. The GSW cryptosystem [183] was the first to propose bootstrap, enabling homomorphic multiplication to transition from a finite number of times to an infinite number, achieving full homomorphic encryption in the truest sense. The CKKS cryptosystem [184], built upon the GSW cryptosystem, enables direct encryption of real number vectors, making it more applicable to machine learning scenarios.
However, different clients may hold different keys, and in Homomorphic Encryption (HE), the same key must be used for encryption to perform ciphertext computation. This poses a challenge for HE in heterogeneous scenarios in federated learning. An effective solution is key distribution, where all client keys are managed and distributed by a trustworthy third party (TTP). This scheme is easy to implement, but the security of the system relies on whether the TTP is truly trustworthy. Another method is called key agreement, which does not require a TTP. Each client communicates on their own, negotiating with the keys held by themselves and the other party to generate a consistent session key. Currently, there is a framework called xMK-CKKS [185], designed based on the key agreement scheme, that can merge all client public keys into one key for CKKS encryption.
IV-B2 Functional Encryption
In conventional cryptographic schemes, possession of the private key implies the ability to fully decrypt the ciphertext. However, functional encryption (FE) [186] imposes a limit on the amount of information that a key can access. In FE, a key $\mathsf{sk}_{f}$ corresponds to a function $f$. Assuming there is a plaintext $x$, an entity possessing $\mathsf{sk}_{f}$ can only obtain $f(x)$ after decryption and cannot know the information of $x$ (unless $f(x)=x$ happens to be true). In a heterogeneous environment, FE requires a TTP to distribute various private keys $\mathsf{sk}$, which implies that the security of the entire system depends on the trustworthiness of the TTP.
In federated learning, the function $f$ corresponding to the key is often set as Multi-Input Functional Encryption [187], as in HybridAlpha [188]. This setting can effectively prevent inference attacks. In response to the security issues of the TTP, Chotard et al. [189] proposed a decentralized scheme, DMCFE, which allows the system to operate without any TTP, thereby enhancing security. Building on this, 2DMCFE [190] combines the advantages of previous research and can effectively resist mix-and-match attacks.
IV-B3 Secrecy Sharing
Secret sharing (SS) [191, 192] is a technique for splitting and storing secret information, which can effectively prevent the leakage of secrets due to collusion among multiple participants. The core idea is to split the secret into $n$ parts, and at least $t$ slices are needed to fully restore the secret. This scheme is called $(t,n)$-secret sharing.
The most classic in SS is the Shamir scheme [191]. In this scheme, the secret is encoded as a $(t-1)$-dimensional polynomial function $P(x)$, and each participant is assigned the information of a point on $P(x)$. In this way, $t$ participants can restore $P(x)$ from the known information of $t$ points according to the Lagrange interpolation method, thereby restoring the secret. The role of the Shamir scheme in federated learning has been discovered by some researchers, such as the SecAgg [193] and SecAgg+ [194] schemes. The SAFELearn [195] scheme and others extend SS to a multi-server situation. FedShare [196], while applying multi-server SS, also optimizes for utility loss and communication costs.
IV-C Model Tracing
In contrast to the methods previously introduced, the methods discussed in this section primarily serve the function of post-hoc tracing. That is, once an attack has been detected, these methods allow for the identification and verification of the source of the attack. While these methods cannot prevent an attack from occurring in advance, they can obtain evidence after the fact, providing administrators or users of the federated learning system with an advantage at the level of rules or law.
IV-C1 Digital watermarking
Digital watermarking [197] is a technique that uses algorithms to add identifiable symbols to target data without affecting the usability of the original information, effectively tracing the source of the data. In the context of federated learning, digital watermarks can be added to the transmitted model information. Even if the information is stolen or leaked, the digital watermark can be used to track or forensically investigate this violation. Some work has already combined digital watermarking technology in FL [198], and corresponding improvements have been proposed for issues such as compression [199] or robustness [200].
IV-C2 Blockchains
Blockchain [201] is a distributed ledger that uses cryptography to provide security guarantees, offering immutable information recording capabilities in insecure environments, initially proposed in Bitcoin . Due to its strong security and traceability, blockchain has been extensively applied in areas such as cryptocurrencies and Non-Fungible Tokens (NFTs) [202]. An important derivative application of blockchain is smart contracts, initially proposed by Ethereum. Smart contracts allow the inclusion of code that can be automatically triggered for execution in the blockchain, and the execution of this code is recorded and verified by the blockchain. In the context of federated learning, the role of blockchain is particularly evident due to the following functionalities:
  • Malicious Client Identification. Blockchain can record the information uploaded by each client, and this information is immutable, which can be used for evidence collection after an attack is discovered.
  • Supervision on Server. Blockchain can also supervise the behavior of servers in scenarios where clients do not trust the servers, checking whether each aggregation behavior is malicious and constitutes an attack.
  • Incentive Mechanism. Blockchain can incentivize clients based on the quality of the model they upload, in the form of blockchain cryptocurrencies, which can prevent free-riding attacks, motivate clients to collect better data for training, and ensure a certain level of fairness.
Several researchers have proposed numerous federated learning frameworks that incorporate blockchain. Among them, CREAT [203] uses smart contracts on the blockchain to verify the reliability of model information uploaded by clients. Toyoda and Zhang [204] proposed a blockchain federated learning algorithm with an incentive mechanism, which scores each client through a voting scheme to achieve competition among clients. BFLC [142], based on blockchain, implements a decentralized federated learning committee mechanism that can be used to resist malicious attacks. Beyond theoretical research, open-source frameworks like VeryFL [205] are advancing this field from a practical perspective.
V Future Research Directions
With the rapid development of big data and artificial intelligence technologies, people are paying increasing attention to privacy data containing a large amount of potential value. As a distributed machine learning paradigm that protects privacy, FL is gradually becoming an important tool for processing distributed data. In heterogeneous scenarios, FL faces multiple challenges, including data heterogeneity, model heterogeneity, task heterogeneity, communication heterogeneity, device heterogeneity, and privacy protection defects. Despite the existence of the above-mentioned research methods, there are still some challenges that have not been resolved. In the following, we will summarize some potential research directions as follows:
Heterogeneity: (1) Data Distribution [206]: Data heterogeneity is one of the main issues faced by FL in heterogeneous scenarios. Future research can start from the data itself, focusing on preprocessing data across clients and designing more intelligent data selection and filtering mechanisms to improve data quality and model training efficiency. At the same time, we can consider data generalization theory to build a bridge between the source domain and the target domain, thus alleviating the problem of data heterogeneity. (2) Models [78]: Model heterogeneity often leads to differences in model structure, parameter scale, and so on. Future research can focus on model integration and fusion techniques, unifying the semantic information of models to ensure that multiple models with different structures have the same semantic branches, supporting low-loss splicing and fusion of models with different structures. Additionally, for model heterogeneity issues, we can explore optimization techniques such as adaptive training adjustments and model pruning to improve the training efficiency and performance of the model. (3) Tasks [207]: Future research can focus on the analysis and utilization of task relevance, improving the generalization ability of the model by exploring the inherent connections between tasks. Furthermore, we can explore dynamic adjustment and adaptation techniques for model tasks to meet the needs of different tasks and improve the adaptability of the model. (4) Communication [208]: Federated learning relies on the robustness of communication. Studying adaptive network topologies can help mitigate the impact of communication heterogeneity on the central server and improve the overall performance of FL. Additionally, we can focus on the design of fault-tolerant mechanisms for FL to mitigate the negative impact of clients being offline on the overall system. Strategies such as optimizing data transmission formats, reducing communication rounds, and designing efficient communication scheduling strategies are also viable solutions. (5) Devices [56]: Device heterogeneity is mainly reflected in differences in the types of devices participating in FL, computing capabilities, storage capacity, and other aspects. Studying the partitioning and distribution strategies of model parameters and techniques such as model compression and quantization will be feasible solutions to address device heterogeneity issues.
Federated Fairness: In heterogeneous scenarios, differences in data distribution, computing resources, and other factors among different participants may lead to fairness issues in the FL process [209]. Future research can focus on improving the fairness of FL algorithms to ensure that different participants have fair opportunities and benefits during model training. Additionally, we can establish fairness evaluation indicators and mechanisms to quantitatively evaluate and optimize the fairness of FL algorithms.
Secure and Trustworthy Strategies: Security and trustworthiness are important prerequisites for the widespread application of FL [210]. Future research can focus on the design and implementation of secure and trustworthy strategies, including participant identity verification, communication encryption, and model update verification. Additionally, we can explore secure and trustworthy solutions based on advanced cryptography techniques such as zero-knowledge proofs and homomorphic encryption to improve the overall security performance of FL.
Incentive Mechanisms: In heterogeneous scenarios, due to differences in the interests and contributions of different participants, it is necessary to design reasonable incentive mechanisms to promote their active participation and cooperation [211]. Future research can focus on the design and optimization of incentive mechanisms, including reward allocation strategies and contribution evaluation methods. Additionally, we can explore incentive mechanism designs based on methods such as game theory and auction theory to achieve more fair and effective resource allocation and interest coordination.
Transfer Learning: Transfer learning can transfer knowledge from one domain to another, thereby accelerating model training in new domains [212]. In federated scenarios, data may be generated in real-time. Therefore, studying how to use FL to achieve cross-domain transfer learning has important practical significance for improving the adaptability and performance of FL.
In conclusion, there are numerous potential research directions in FL, especially in heterogeneous scenarios. Future research should focus on addressing these challenges and exploring new techniques to improve the performance, efficiency, and fairness of FL.
VI Conclusion
In this paper, we provide insights into key concepts in the field of heterogeneous federated learning and the multiple challenges it faces. In particular, we analyze in detail the heterogeneity of data, models, tasks, communications, and devices, which pose numerous algorithmic challenges in federated learning systems. To address these challenges, in this paper, we systematically categorize existing research approaches by classifying them into data-level, model-level, and structure-level according to the processing hierarchy, each of which corresponds to a specific problem and solution strategy. Our categorization approach provides a clear perspective for understanding the efficacy of various types of algorithms in coping with heterogeneous environments. With this hierarchical strategy, researchers and practitioners can more efficiently select methods that fit their specific needs. Finally, this paper looks at future research directions in heterogeneous federated learning, which will undoubtedly continue to be a dynamic and challenging research direction in the field of machine learning as technology advances and application needs increase. We look forward to more innovative research results in this field in the future to promote the development of technology and popularization of applications.
Acknowledgment
The research is supported by the National Key Research and Development Program of China (2023YFB2703700), the National Natural Science Foundation of China (62176269), the Guangzhou Science and Technology Program (2023A04J0314), the National Natural Science Foundation of China and Guangdong Provincial Joint Fund (No. U1911202).